Section 16

Applications

In case of several authors, Invited Speakers are marked with a ^{∗}.

Marco Avellaneda: The Minimum-Entropy Algorithm and Related Methods for Calibrating Asset-Pricing Models . . . . . . . . . . . . . . . . . . . . . III 545

Andreas Dress^{∗}, Werner Terhalle: The Tree of Life and Other Affine Buildings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 565

Leslie Greengard^{∗} and Xiaobai Sun: A New Version of the Fast Gauss Transform . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 575

Ulf Grenander: Strategies for Seeing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 585

Frank Hoppensteadt^{∗} and Eugene Izhikevich: Canonical Models in Mathematical Neuroscience . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 593

Thomas Yizhao Hou: Numerical Study of Free Interface Problems Using Boundary Integral Methods . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 601

Gérard Iooss: Travelling Water-Waves, as a Paradigm for

Bifurcations in Reversible Infinite Dimensional “Dynamical” Systems III 611

Yury Grabovsky and Graeme W. Milton^{∗}: Exact Relations for Composites: Towards a Complete Solution . . . . . . . . . . . . . . . . . . . . . . . . . III 623

Charles S. Peskin: Optimal Dynamic Instability of Microtubules . . III 633

Documenta Mathematica · Extra Volume ICM 1998 · III ·

544

Documenta Mathematica · Extra Volume ICM 1998 · III ·